# Spatiotemporal Density Correction of Multivariate Global Climate Model Projections using Deep Learning

### v2

@@ Abstract
Global Climate Models (GCMs) are numerical models that simulate complex physical processes within the Earth’s climate system and are essential for understanding and predicting climate change. However, GCMs suffer from systemic biases due to simplifications made to the underlying physical processes. GCM output therefore needs to be bias corrected before it can be used for future climate projections. Most common bias correction methods, however, cannot preserve spatial, temporal, or inter-variable dependencies. We propose a new semi-parametric conditional density estimation (SPCDE) for density correction of the joint distribution of daily precipitation and maximum temperature data obtained from gridded GCM spatial fields. The Vecchia approximation is employed to preserve dependencies in the observed field during the density correction process, which is carried out using semi-parametric quantile regression. The ability to calibrate joint distributions of GCM projections has potential advantages not only in estimating extremes, but also in better estimating compound hazards, like heat waves and drought, under potential climate change. Illustration on historical data from 1951–2014 over two $5\times 5$ spatial grids in the US indicate that SPCDE can preserve key marginal and joint distribution properties of precipitation and maximum temperature, and predictions obtained using SPCDE are better calibrated compared to predictions using asynchronous quantile mapping and canonical correlation analysis, two commonly used bias correction approaches.
@@ 1 Introduction
Global Climate Models (GCMs) are numerical models to explain complex interactions between the ocean, atmosphere, and land surface. GCMs are used to study long-term climate trends and to provide projections based on possible future scenarios associated with, e.g., greenhouse gas and aerosol emissions, land use, and population growth [^1] https://www.climate.gov/maps-data/climate-data-primer/predicting-climate/climate-models. The latest generation of GCMs are collectively known as the Climate Model Intercomparison Project version 6 (CMIP6; Eyring et al., 2016). It contains the outputs, primarily ocean, atmospheric, and land surface conditions, as ensembles at a coarser spatial scale for further analyzing impacts on various sectors. For instance, GCM forcings, typically precipitation and temperature, are used to assess for both surface water and groundwater availability over the planning period (Singh et al., 2015; Seo et al., 2016). To develop these assessments of future water availability, GCM projections, particularly for precipitation and temperature, are forced through hydrologic models, which are subsequently used as an input into a reservoir model (Singh et al., 2015; Seo et al., 2016). Projected inflow into reservoirs are typically used for allocating water for multipurpose uses that includes municipal, irrigation, hydropower and flood control. GCM projections for other variables, particularly temperature and humidity, are also critical for quantifying energy demand (Eshraghi et al., 2021).
@@
Traditional bias correction methods like quantile mapping (QM) have been used to adjust the distribution of GCM data by mapping quantiles from the GCM to the corresponding quantiles in the observational record (Stoner et al., 2013; Thrasher et al., 2012; Maraun, 2013). While effective for single-variable adjustments, QM often fails to account for inter-variable dependencies (Bhowmik and Sankarasubramanian, 2019). Multivariate bias correction methods aim to preserve the relationships between climatic variables (Bhowmik et al., 2017). While multivariate QMs attempt to address this issue (Piani and Haerter, 2012; Cannon, 2018), QM is inappropriate for the bias correction of future projections as it distorts future climate change in a nonphysical way (e.g., Maurer and Pierce, 2014). Other commonly used methods such as bias corrected spatial downscaling (BCSD, Wood et al., 2002), bias corrected constructed analogues (BCCA, Maurer et al., 2010), and multivariate adapted constructed analogues (MACA, Abatzoglou and Brown, 2012) also employ QM. Multivariate methods like localized constructed analogs (LOCA; Pierce et al., 2014, 2023) and asynchronous canonical correlation analysis (CCA; Bhowmik et al., 2017) offer improvements by better capturing the temporal covariability in climate variables (e.g., cross-correlation), but do not explicitly consider spatial dependence during bias correction. Finally, recent deep learning-based methods, like DeepSD (Vandal et al., 2017), leverage the power of deep learning algorithms to model complex, nonlinear relationships between climate variables and local physiographical information such as elevation, but have not been evaluated for spatiotemporal dependency in multiple climate variables. Additionally, none of the aforementioned methods, with the exception of LOCA v2 (Pierce et al., 2023), are geared towards reproducing observed extreme quantiles of climate variables. However, LOCA v2 also fails to preserve spatial dependency. We argue that bias correction should be viewed as a problem of density correction - calibrating the spatiotemporal joint density of the historical GCM projections to the spatiotemporal joint density of the observed multivariate climate data. Subsequently, the ‘calibration’ model will be run with future GCM projections to develop density-corrected (i.e., calibrated) GCM projections that remove the systematic biases in the GCM joint densities. Our effort here is to develop such a density correction approach by comparing the joint density fields of GCM projections with its observed counterpart.
[…]
@@
The proposed SPCDE connects conditional density estimation (CDE; Hyndman et al., 1996; Li and Racine, 2006) and bias correction in a unique manner, and in the process mitigates several drawbacks of commonly used bias correction methods. A key aspect of SPCDE (described in further detail in Sections 2.32.5) is the estimation of the joint distribution of GCM and observational data within a single SPQR model, and thereafter using cumulative density function (CDF) and quantile function (QF) transforms to calibrate the model data. Similar CDF based approaches have been used for bias correcting future climate projections (e.g., Michelangeli et al., 2009; Li et al., 2010); however, these approaches estimate the CDF of the GCM and observed data independent of each other. SPCDE also bears resemblance to normalizing flows (NFs; Rezende and Mohamed, 2015; Papamakarios et al., 2021; Kobyzev et al., 2021) and transport maps (Marzouk et al., 2016; Katzfuss and Schäfer, 2024). Both of these methods aim to transform a (usually complex) target distribution to a much simpler reference distribution (e.g., a standard Gaussian) through a series of invertible transformations; the transformations are referred to as flows and transport maps, respectively. Inverting the transformations allows for sampling from the target distribution as well as making predictions. Transport maps can be considered a special case of an NF (Katzfuss and Schäfer, 2024). Similarly, SPCDE parallels NFs - we can consider the GCM and observational data to be distributed according to the reference and target distributions respectively, with the flows being a sequence of CDFs and QFs. This also highlights the key difference between SPCDE and NFs. While the reference distribution in NFs is assumed to be known, the joint distribution of the GCM data also needs to be learned alongside the joint distribution of the observational data. While this is a more complex task, SPCDE’s ability to estimate them simultaneously makes the calibration step of density correction quite straightforward.
@@
The remainder of this paper is organized as follows. Section 2 introduces the data and methods for our study. Section 3 contains a case study for density correction of climate model data at two different regions of the US. Section 4 concludes.
@@ 2.1 Overview
Section 2.2 describes the different datasets used in our study, including resolution and pre-processing used to align the different datasets on a common spatial grid. Section 2.3 introduces the SPCDE approach for density correction of TMAX and PRCP at a single grid cell. Sections 2.4 and 2.5 extend SPCDE to spatial and spatiotemporal settings, respectively. Section 2.6 outlines SPQR, the conditional density estimator that is used for SPCDE. Finally, Section 2.7 describes the metrics used to assess the quality of the calibrated data.
@@ 2.2 Temperature and precipitation datasets
[…]
@@ 2.3 Density correction model for GCM projections at a single location
Our proposed density correction model has three steps - estimation, projection, and calibration. We begin by stating our model for the general multivariate case in this section; explicit formulations for spatial and spatiotemporal data are described in subsequent sections. We build a statistical model for $p$ continuous responses, $\mbox{\bf Y}=(Y_{1},...,Y_{p})$ at a single location, which are assumed to have invertible marginal quantile and distribution functions that depends on variables Z (e.g., season or elevation). The model is described simultaneously for both climate model output and observational data, with an auxiliary variable $X=0$ indicating the observation is from the climate model and $X=1$ indicating it is from the observational record.
@@
where $\mbox{\bf Y}_{(j)}=\{Y_{i}:i\in\mathcal{N}_{j}\}$ and the conditioning set is $\mathcal{N}_{j}=\{1,...,j-1\}$, which depends on the ordering of the $p$ variables. The responses Y can alternatively be represented as transformations of the latent variables $\mbox{\bf U}=(U_{1},...,U_{p})$, with $U_{j}\stackrel{{\scriptstyle iid}}{{\sim}}\mbox{Uniform}(0,1)$ for $j\in\{1,...,p\}$. The map from U to Y can be expressed sequentially,
@@
While (4) holds for any ordering of the $p$ variables, certain orderings will likely to lead to better estimation depending on the complexity of the $p$ densities. In the simple case of TMAX and PRCP at a single location (i.e., $p=2$), a natural ordering would be based on choosing $Y_{1}$ as TMAX and $Y_{2}$ as PRCP. This decomposes the conditional distribution of PRCP and TMAX as a conditional and a marginal distribution, assuming PRCP depends on TMAX. Physically, TMAX and PRCP are coupled, and choosing $Y_{1}$ as PRCP and $Y_{2}$ as TMAX also provides a valid statistical formulation. However, we assume that PRCP depends on TMAX for somewhat practical reasons. Intuitively, the distribution of TMAX has a much simpler functional form compared to PRCP. PRCP has a preponderance of zeros, and is therefore harder to estimate compared to TMAX. Our proposed dependence structure mitigates this by using TMAX as a predictor for PRCP. Our formulation is also validated in practice; SPCDE where TMAX depended on PRCP led to noticeably worse predictions over both the training and testing periods.
@@
     Estimate $F_{1}$ and $Q_{1}$ for $Y_{1}|\mbox{\bf Z},X$triangleright Estimation step
     Estimate $F_{2}$ and $Q_{2}$ for $Y_{2}|Y_{1},\mbox{\bf Z},X$triangleright Estimation step
@@
     $u_{1}\leftarrow F_{1}(Y_{1}|X=0)$triangleright Projection step
     $Y_{1}^{*}\leftarrow Q_{1}(u_{1}|X=1)$triangleright Calibration step
@@
     $\mbox{\bf Y}_{(2)}\leftarrow Y_{1}$
     $u_{2}\leftarrow F_{2}(Y_{2}|\mbox{\bf Y}_{(2)},X=0)$triangleright Projection step
@@
     $Y_{2}^{*}\leftarrow Q_{2}(u_{2}|\mbox{\bf Y}_{(2)}^{*},X=1)$triangleright Calibration step
@@ 2.4 Extension to spatial data
where $\mbox{\bf Y}_{(j)}=\{Y_{i};i\in\mathcal{N}_{j}\}$ and the conditioning set is $\mathcal{N}_{j}\subseteq\{1,...,j-1\}$. We refer to $\mbox{\bf Y}_{(j)}$ as the Vecchia neighbor set going forward. The size of the conditioning set $\mathcal{N}_{j}$ is therefore bounded above, i.e., $|\mathcal{N}_{j}|\leq j-1$. In practice, $10\leq|\mathcal{N}_{j}|\leq 30$ are often sufficient even when $p$ is very large; we refer the reader to Guinness (2018) and Katzfuss and Guinness (2021) for discussions on optimal choices for cardinality and members of $\mathcal{N}_{j}$. The Vecchia approximation facilitates the decomposition of a complex, high-dimensional response vector into a product of independent univariate likelihoods. It is important to note that the approximation is applied only to the dependence structure within Y, and the right hand expression of (5) is still a valid joint distribution for Y.
@@
To apply the Vecchia decomposition to the joint distribution of TMAX and PRCP at $n$ locations (with the extension to more than two variables being straightforward), let $Y_{1l}$ and $Y_{2l}$ denote the response for TMAX and PRCP at location $l\in\{1,\ldots,n\}$ corresponding to grid point $\mbox{\bf s}_{l}=(s_{l1},s_{l2})$. The $p=2n$ variables are therefore ordered as $\mbox{\bf Y}=(Y_{11},Y_{21},\ldots,Y_{1n},Y_{2n})$. In this study, we choose the Vecchia neighboring set for $Y_{1l}$ to be $\mbox{\bf Y}_{{(1l)}}=\{Y_{1,l-1},\ldots,Y_{1,l-m}\}$, and the Vecchia neighboring set for $Y_{2l}$ as $\mbox{\bf Y}_{{(2l)}}=\{Y_{1l},\ldots,Y_{1,l-m},Y_{2,l-1},\ldots,Y_{2,l-m}\}$. This encapsulates our prior assumption that PRCP depends on TMAX but not vice versa. Therefore, while TMAX at a location depends on TMAX at its Vecchia neighbor locations, PRCP depends on TMAX at the current location and the PRCP at the (same) Vecchia neighbor locations. While other formulations are possible, we found this to be intuitive and tractable.
2.5 Extension to spatiotemporal data
Algorithm 2 Spatiotemporal density correction for TMAX and PRCP
Data $\mbox{\bf Y},X$
for $l=1,\ldots,n$ do
     $\mbox{\bf Y}_{(1lt)}=\{Y_{1l,t-1},Y_{1,l-1,t},\ldots,Y_{1,l-m,t}\}$
     procedure SPQR for TMAX and PRCP
         Estimate $F_{1l}$ and $Q_{1l}$ for $Y_{1lt}|\mbox{\bf Y}_{(1lt)},X$triangleright Estimation step
         Estimate $F_{2l}$ and $Q_{2l}$ for $Y_{2lt}|\mbox{\bf Y}_{(2lt)},X$triangleright Estimation step
     end procedure
     procedure TMAX density correction
         $u_{1lt}\leftarrow F_{1l}(Y_{1lt}|\mbox{\bf Y}_{(1lt)},X=0)$triangleright Projection step
         for $t=2,\ldots,T$ do
              $Y_{1lt}^{*}\leftarrow Q_{1l}(u_{1lt}|\mbox{\bf Y}_{(1lt)}^{*},X=1)$triangleright Calibration step
         end for
     end procedure
     procedure PRCP density correction
         $u_{2lt}\leftarrow F_{2l}(Y_{2lt}|\mbox{\bf Y}_{(2lt)},X=0)$triangleright Projection step
         for t = 2,…,T do
              $Y_{2lt}^{*}\leftarrow Q_{2l}(u_{2lt}|\mbox{\bf Y}_{(2lt)}^{*},X=1)$triangleright Calibration step
         end for
     end procedure
end for
Finally, we extend the model from Section 2.4 to incorporate temporal information. Let $Y_{1lt}$ and $Y_{2lt}$ denote the response for TMAX and PRCP at location $l\in\{1,\ldots,n\}$ and time $t\in\{1,...,T\}$. We order the $p=2nT$ variables as $(Y_{111},\ldots,Y_{11T},Y_{211},\ldots,Y_{21T},\ldots,Y_{1n1},\ldots,Y_{1nT},Y_{2n1},\ldots,Y_{2nT})$. The model in Section 2.4 applies to multivariate time series data after key assumptions of Markovian and stationarity dependence structure. The Markovian structure states that the conditional quantile function for $Y_{jlt},j=1,2,$ given all observation before time $t$ depends only on data from the previous $T^{\prime}<T$ time steps; the stationarity assumption states that the conditional quantile functions are invariant to $t$.
In this study, we assume the conditional quantile function to be first order Markov in time, i.e., $T^{\prime}=1$. This assumption is reasonable given the selected are from the Southeast and Southwest US in regions where pronounced seasonality in precipitation is absent (Petersen et al., 2012). This implies that the Vecchia neighbor set for $Y_{1lt}$ is $\mbox{\bf Y}_{(1lt)}=\{Y_{1l,t-1},Y_{1,l-1,t},\ldots,Y_{11t}\}$, and for $Y_{2lt}$ is $\mbox{\bf Y}_{(2lt)}=\{Y_{2l,t-1},Y_{1l,t-1},Y_{1,l,t},\ldots,\\ Y_{1,l-m,t},Y_{2,l-1,t},\ldots,Y_{2,l-m,t}\}$. These are straightforward extensions of the Vecchia neighbor sets presented in Section 2.4, with TMAX on day $t$ being additionally depending on TMAX on day $t-1$, and PRCP on day $t$ additionally depending on both PRCP and TMAX from day $t-1$. The conditional QF $Q$, CDF $F$, and calibration function $C$, for $\mbox{\bf Y}_{lt}$ are defined identically to Section 2.3; these do not depend on $t$ due to the stationarity assumption. Algorithm 2 details the density correction methodology for spatiotemporal data.

### v3

# Density correction for multivariate spatial fields of global climate model output using deep learning

@@ Abstract
Global Climate Models (GCMs) are numerical models that simulate complex physical processes within the Earth’s climate system and are essential for understanding and predicting climate change. However, GCMs suffer from systemic biases due to simplifications made to the underlying physical processes. GCM output therefore needs to be bias corrected before it can be used for future climate projections. Most common bias correction methods, however, cannot preserve spatial, temporal, or inter-variable dependencies. We propose a new semi-parametric estimation of conditional densities (SPECD) approach for density correction of the joint distribution of daily precipitation and maximum temperature data obtained from gridded GCM spatial fields. The Vecchia approximation is employed to preserve dependencies in the observed field during the density correction process, which is carried out using semi-parametric quantile regression. The ability to calibrate joint distributions of GCM projections has potential advantages not only in estimating extremes, but also in better estimating compound hazards, like heat waves and drought, under potential climate change. Illustration on historical data from 1951–2014 over two $5\times 5$ spatial grids in the US indicate that SPECD can preserve key marginal and joint distribution properties of precipitation and maximum temperature, and predictions obtained using SPECD are better calibrated compared to predictions using asynchronous quantile mapping and canonical correlation analysis, two commonly used bias correction approaches.
@@ 1 Introduction
Global Climate Models (GCMs) are numerical models to explain complex interactions between the ocean, atmosphere, and land surface. GCMs are used to study long-term climate trends and to provide projections based on possible future scenarios associated with, e.g., greenhouse gas and aerosol emissions, land use, and population growth (Gettelman and Rood, 2016). The latest generation of GCMs are collectively known as the Climate Model Intercomparison Project version 6 (CMIP6; Eyring et al., 2016). It contains the outputs, primarily ocean, atmospheric, and land surface conditions, as ensembles at a coarser spatial scale for further analyzing impacts on various sectors. For instance, GCM forcings (typically precipitation and temperature) are used to assess for both surface water and groundwater availability over planning periods (Singh et al., 2015; Seo et al., 2016). To develop these assessments of future water availability, GCM projections, particularly for precipitation and temperature, are forced through hydrologic models, which are subsequently used as an input into a reservoir model (Singh et al., 2015; Seo et al., 2016). Projected inflow into reservoirs are typically used for allocating water for multipurpose uses that includes municipal, irrigation, hydropower and flood control. GCM projections for other variables, particularly temperature and humidity, are also critical for quantifying energy demand (Eshraghi et al., 2021).
@@
Traditional bias correction methods like asynchronous quantile mapping (QM) have been used to adjust the distribution of GCM data by mapping quantiles from the GCM to the corresponding quantiles in the observational record (Stoner et al., 2013; Thrasher et al., 2012; Maraun, 2013). While effective for single-variable adjustments, QM often fails to account for inter-variable dependencies (Bhowmik and Sankarasubramanian, 2019). Multivariate bias correction methods aim to preserve the relationships between climatic variables (Bhowmik et al., 2017). While multivariate QMs attempt to address this issue (Piani and Haerter, 2012; Cannon, 2018), QM is inappropriate for the bias correction of future projections as it distorts future climate change in a nonphysical way (e.g., Maurer and Pierce, 2014). Other commonly used methods such as bias corrected spatial downscaling (BCSD, Wood et al., 2002), bias corrected constructed analogues (BCCA, Maurer et al., 2010), and multivariate adapted constructed analogues (MACA, Abatzoglou and Brown, 2012) also employ variants of QM as part of their framework. Multivariate methods like localized constructed analogs (LOCA; Pierce et al., 2014, 2023) and asynchronous canonical correlation analysis (CCA; Bhowmik et al., 2017) offer improvements by better capturing the temporal covariability in climate variables (e.g., cross-correlation), but do not explicitly consider spatial dependence during bias correction. Finally, recent deep learning-based methods, like DeepSD (Vandal et al., 2017), leverage the power of deep learning algorithms to model complex, nonlinear relationships between climate variables and local physiographical information such as elevation, but have not been evaluated for the dependency between multiple climate variables, or the spatial dependence within each variable. Additionally, none of the aforementioned methods, with the exception of LOCA v2 (Pierce et al., 2023), are geared towards reproducing observed extreme quantiles of climate variables. However, LOCA v2 also fails to preserve spatial dependency. We argue that bias correction should be viewed as a problem of density correction - calibrating the spatiotemporal joint density of the historical GCM projections to the spatiotemporal joint density of the observed multivariate climate data. Subsequently, the ‘calibration’ model will be run with future GCM projections to develop density-corrected (i.e., calibrated) GCM projections that remove the systematic biases in the GCM joint densities. Our effort here is to develop such a density correction approach by comparing the joint density fields of GCM output with its observed counterpart.
In this paper, we develop a new semi-parametric estimation of conditional densities (SPECD) approach for density correction of the joint distribution of daily precipitation (PRCP) and maximum temperature (TMAX) data. The joint distribution of the two variables over a spatial field is decomposed into a product of univariate conditional distributions using a Vecchia approximation (Vecchia, 1988), and each univariate density is estimated using semi-parametric quantile regression (SPQR; Xu and Reich, 2021). SPQR uses neural networks to learn univariate conditional densities of random variables. The deep learning model underpinning SPQR ensures that nonlinear relationships are captured, and the Vecchia approximation ensures that spatial and cross correlations are preserved. We test our method with data over twenty five 100 km×100 $\mbox{ km}\times 100\mbox${ km} grid cells in the Southeast (SE) and Southwest (SW) US, depicted in Figure 1. We use daily data from 1951–2000 to train our density correction model, and data from 2001–2014 to validate our results. Our results indicate that SPECD is able to preserve key marginal and joint distribution properties of TMAX and PRCP. Additionally, SPECD predictions are better calibrated over training and validation periods compared to QM and CCA for our spatial domain.
@@
The proposed SPECD connects conditional density estimation (CDE; Hyndman et al., 1996; Li and Racine, 2006) and bias correction in a unique manner, and in the process mitigates several drawbacks of commonly used bias correction methods. A key aspect of SPECD (described in further detail in Section 2.3) is the estimation of the joint distribution of GCM and observational data within a single SPQR model, and thereafter using cumulative density function (CDF) and quantile function (QF) transforms to calibrate the model data. Similar CDF based approaches have been used for bias correcting future climate projections (e.g., Michelangeli et al., 2009; Li et al., 2010); however, these approaches estimate the CDF of the GCM and observed data independent of each other. SPECD also bears resemblance to normalizing flows (NFs; Rezende and Mohamed, 2015; Papamakarios et al., 2021; Kobyzev et al., 2021) and transport maps (Marzouk et al., 2016; Katzfuss and Schäfer, 2024). Both of these methods aim to transform a (usually complex) target distribution to a much simpler reference distribution (e.g., a standard Gaussian) through a series of invertible transformations; the transformations are referred to as flows and transport maps, respectively. Inverting the transformations allows for sampling from the target distribution as well as making predictions. Transport maps can be considered a special case of an NF (Katzfuss and Schäfer, 2024). Similarly, SPECD parallels NFs - we can consider the GCM and observational data to be distributed according to the reference and target distributions respectively, with the flows being a sequence of CDFs and QFs. This also highlights the key difference between SPECD and NFs. While the reference distribution in NFs is assumed to be known, the joint distribution of the GCM data also needs to be learned alongside the joint distribution of the observational data. While this is a more complex task, SPECD’s ability to estimate them simultaneously makes the calibration step of density correction quite straightforward.
@@
The remainder of this paper is organized as follows. Section 2 introduces the data and methods for our study. Section 3 studies the performance of SPECD using a numerical study. Section 4 contains a case study for density correction of climate model data at two different regions of the US. Section 5 concludes.
@@ 2.1 Overview
Section 2.2 describes the different datasets used in our study, including resolution and pre-processing used to align the different datasets on a common spatial grid. Section 2.3 introduces the SPECD approach for density correction of TMAX (Kelvin) and PRCP (mm) at a single grid cell. Sections 2.4 extends SPECD to the spatial setting. Section 2.5 outlines SPQR, the conditional density estimator that is used for SPECD. Finally, Section 2.6 describes the metrics used to assess the quality of the calibrated data.
@@ 2.2 Temperature and precipitation datasets
Figure 2: Monthly mean fields of nClimGrid anomalies (top) and GCM raw data (bottom) of TMAX and PRCP for the Southeast US study region.
The study utilizes an ensemble member from the Geophysical Fluid Dynamics Laboratory (GFDL) GCM of of the CMIP6 experiments. The model, developed by NOAA, provides data at a spatial resolution of $100\,\mbox{km}\times 100\,\mbox{km}$, which are used to analyze and density-correct the joint distribution of daily TMAX and PRCP for selected grids in the Southeast US and the Southwest US which have very distinct climatology. The Southwest grid (Figure 1, left panel) in the four corners region is arid and has complex topography, spanning the states of Utah, Colorado, Arizona, and Mexico. The Southeast grid (Figure 1, right panel) is humid, and spans the states of Indiana, Ohio, West Virginia, Kentucky, Virginia, Tennessee, North Carolina, South Carolina, Alabama, and Georgia. The GCM data are obtained from Earth System Grid Federation (ESGF) World Climate Research Programme (WCRP). The GFDL model is meant to be representative of CMIP6, and our methodology can be applied to any GCM in principle. The primary observational data comes from NOAA’s nClimGrid-daily data product (Vose et al., 2014; Durre et al., 2022), which offers daily precipitation and temperature data from 1951 to the present at a 5 km×5 $\mbox{ km}\times 5\mbox${ km} resolution. The observed climate data is aggregated to the same resolution as the GCM data, such that all the datasets used in our study are spatially aligned. The aggregation is carried out by taking the mean of the nClimGrid pixels using the Climate Data Operators software (CDO; Schulzweida, 2023). Figure 2 plots the monthly mean fields for July 2014 TMAX and PRCP in the Southeast region. The top row shows nClimGrid anomalies, calculated by subtracting the long-term mean of the data, at its native resolution. The bottom row shows the raw GCM data for the same time period. We use 1951–2014 as our study period, since the CMIP6 historical simulations run until 2014, and the observational record begins at 1950.
@@ 2.3 Density correction model for GCM projections at a single location
Our proposed density correction model has three steps - estimation, projection, and calibration. We begin by stating our model for the general multivariate case in this section; explicit formulations for spatial and spatiotemporal data are described subsequently. We build a statistical model for $p$ continuous responses, $\mbox{\bf Y}=(Y_{1},...,Y_{p})$ at a single location, which are assumed to have invertible marginal quantile and distribution functions that depends on variables Z (e.g., season or elevation). The model is described simultaneously for both climate model output and observational data, with an auxiliary variable $X=0$ indicating the observation is from the climate model and $X=1$ indicating it is from the observational record.
@@
where $\mbox{\bf Y}_{(j)}=\{Y_{i}:i\in\mathcal{N}_{j}^{\prime}\}$ and the conditioning set is $\mathcal{N}_{j}^{\prime}=\{1,...,j-1\}$, which depends on the ordering of the $p$ variables. The responses Y can alternatively be represented as transformations of the latent variables $\mbox{\bf U}=(U_{1},...,U_{p})$, with $U_{j}\stackrel{{\scriptstyle iid}}{{\sim}}\mbox{Uniform}(0,1)$ for $j\in\{1,...,p\}$. The map from U to Y can be expressed sequentially,
@@
  Estimate $F_{1}$ and $Q_{1}$ for $Y_{1}|\mbox{\bf Z},X$ $\triangleright$ Estimation step
  Estimate $F_{2}$ and $Q_{2}$ for $Y_{2}|\mbox{\bf Y}_{(2)},\mbox{\bf Z},X$ $\triangleright$ Estimation step
@@
  $u_{1}\leftarrow F_{1}(Y_{1}|\mbox{\bf Z},X=0)$ $\triangleright$ Projection step
  $Y_{1}^{*}\leftarrow Q_{1}(u_{1}|\mbox{\bf Z},X=1)$ $\triangleright$ Calibration step
@@
  $u_{2}\leftarrow F_{2}(Y_{2}|\mbox{\bf Y}_{(2)},\mbox{\bf Z},X=0)$ $\triangleright$ Projection step
@@
  $Y_{2}^{*}\leftarrow Q_{2}(u_{2}|\mbox{\bf Y}_{(2)}^{*},\mbox{\bf Z},X=1)$ $\triangleright$ Calibration step
@@
While (4) holds for any ordering of the $p$ variables, certain orderings will likely to lead to better estimation depending on the complexity of the $p$ densities. In the simple case of TMAX and PRCP at a single location (i.e., $p=2$), a natural ordering would be based on choosing $Y_{1}$ as TMAX and $Y_{2}$ as PRCP. This decomposes the conditional distribution of PRCP and TMAX as a conditional and a marginal distribution, assuming PRCP depends on TMAX. Physically, TMAX and PRCP are coupled, and choosing $Y_{1}$ as PRCP and $Y_{2}$ as TMAX also provides a valid statistical formulation. However, we assume that PRCP depends on TMAX for somewhat practical reasons. Intuitively, the distribution of TMAX has a much simpler functional form compared to PRCP. PRCP has a preponderance of zeros, and is therefore harder to estimate compared to TMAX. Our proposed dependence structure mitigates this by using TMAX as a predictor for PRCP. Our formulation is also validated in practice; SPECD where TMAX depended on PRCP led to noticeably worse predictions over both the training and validation periods.
@@ 2.4 Extension to spatial data
Algorithm 2 Density correction for $Y_{1}=$ TMAX and $Y_{2}=$ PRCP at multiple locations
Data $\mbox{\bf Y},X$
for $l=1,\ldots,n$ do
  $\mbox{\bf Y}_{(1l)}=\{Y_{1,l-1},\ldots,Y_{1,l-m}\}$
  $\mbox{\bf Y}_{(2l)}=\{Y_{1,l},\ldots,Y_{1,l-m},Y_{2,l-1},\ldots,Y_{2,l-m}\}$
  procedure SPQR for TMAX and PRCP
   Estimate $F_{1l}$ and $Q_{1l}$ for $Y_{1l}|\mbox{\bf Y}_{(1l)},X$ $\triangleright$ Estimation step
   Estimate $F_{2l}$ and $Q_{2l}$ for $Y_{2l}|\mbox{\bf Y}_{(2l)},X$ $\triangleright$ Estimation step
  end procedure
  procedure TMAX density correction
   $u_{1l}\leftarrow F_{1l}(Y_{1l}|\mbox{\bf Y}_{(1l)},X=0)$ $\triangleright$ Projection step
   $\mbox{\bf Y}_{(1l)}^{*}=\{Y_{1,l-1}^{*},\ldots,Y_{1,l-m}^{*}\}$
   $Y_{1l}^{*}\leftarrow Q_{1l}(u_{1l}|\mbox{\bf Y}_{(1l)}^{*},X=1)$ $\triangleright$ Calibration step
  end procedure
  procedure PRCP density correction
   $u_{2l}\leftarrow F_{2l}(Y_{2l}|\mbox{\bf Y}_{(2l)},X=0)$ $\triangleright$ Projection step
   $\mbox{\bf Y}_{(2l)}^{*}=\{Y_{1,l}^{*},\ldots,Y_{1,l-m}^{*},Y_{2,l-1}^{*},\ldots,Y_{2,l-m}^{*}\}$
   $Y_{2l}^{*}\leftarrow Q_{2l}(u_{2l}|\mbox{\bf Y}_{(2l)}^{*},X=1)$ $\triangleright$ Calibration step
  end procedure
end for
@@
where $\mbox{\bf Y}_{(j)}=\{Y_{i};i\in\mathcal{N}_{j}\}$ and the conditioning set is $\mathcal{N}_{j}\subseteq\mathcal{N}_{j}^{\prime}=\{1,...,j-1\}$. We refer to $\mbox{\bf Y}_{(j)}$ as the Vecchia neighbor set going forward. The size of the conditioning set $\mathcal{N}_{j}$ is therefore bounded above, i.e., $|\mathcal{N}_{j}|\leq j-1$. In practice, $10\leq|\mathcal{N}_{j}|\leq 30$ are often sufficient even when $p$ is very large; we refer the reader to Guinness (2018) and Katzfuss and Guinness (2021) for discussions on optimal choices for cardinality and members of $\mathcal{N}_{j}$. The Vecchia approximation facilitates the decomposition of a complex, high-dimensional response vector into a product of independent univariate likelihoods. It is important to note that the approximation is applied only to the dependence structure within Y, and the right hand expression of (5) is still a valid joint distribution for Y.
@@
To apply the Vecchia decomposition to the joint distribution of TMAX and PRCP at $n$ locations (with the extension to more than two variables being straightforward), let $Y_{1l}$ and $Y_{2l}$ denote the response for TMAX and PRCP at location $l\in\{1,\ldots,n\}$ corresponding to grid point $\mbox{\bf s}_{l}=(s_{l1},s_{l2})$. The $p=2n$ variables are therefore ordered as $\mbox{\bf Y}=(Y_{11},Y_{21},\ldots,Y_{1n},Y_{2n})$. In this study, we choose the Vecchia neighboring set for $Y_{1l}$ to be $\mbox{\bf Y}_{{(1l)}}=\{Y_{1,l-1},\ldots,Y_{1,l-m}\}$, and the Vecchia neighboring set for $Y_{2l}$ as $\mbox{\bf Y}_{{(2l)}}=\{Y_{1l},\ldots,Y_{1,l-m},Y_{2,l-1},\ldots,Y_{2,l-m}\}$. This encapsulates our prior assumption that PRCP depends on TMAX but not vice versa. Therefore, while TMAX at a location depends on TMAX at its Vecchia neighbor locations, PRCP depends on TMAX at the current location and the PRCP at the (same) Vecchia neighbor locations. While other formulations are possible, we found this to be intuitive and tractable. Algorithm 2 summarizes the calibration process for TMAX and PRCP at multiple locations.
Note that it is possible to extend the ideas of the spatial setting into a spatiotemporal version of the algorithm, where time-lagged variables are used as covariates to capture autocorrelation. Appendix A outlines this extension to the spatiotemporal setting.